### v1

@@ 5. Euler characteristics of perverse sheaves on ${\mathscr{A}}_{g,n}$
6. Key Inequality
In this section, we shall work in the setting of varieties over an algebraically closed field $k$ of characteristic 00. Given a variety $X$, we say that $X$ has the perverse Euler characteristic, or $PEC$-property, or simply that $X$ is $PEC$, if for every perverse sheaf ${\mathcal{K}}\in\operatorname{Perv}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. If $X$ comes equipped with a ${\mathbb{G}}_{m}$-action, we will use similar terminology in the setting of monodromic perverse sheaves: $X$ has the $PEC_{m}$ property if for all ${\mathcal{K}}\in Perv_{m}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. We shall say that $X$ has the $PEC[d]$ (resp. $PEC_{m}[d]$)-property if $\chi(X,{\mathcal{K}}[d])\geq 0$ for all ${\mathcal{K}}\in\operatorname{Perv}(X)$ (resp. $\operatorname{Perv}_{m}(X))$. Clearly $PEC[d]$ is the same as $PEC$ if $d$ is even, and $PEC[-1]$ otherwise.
@@ 6. Key Inequality
Let $(X,Y,f,U,Z)$ be an admissible datum as above. Suppose that $Y$ and $Z$ have the $PEC[d]$-property. Then $X$ also has the $PEC[d]$-property. The analogous assertion holds in the monodromic setting, i.e. when the datum admits a ${\mathbb{G}}_{m}$-action and $Y$ and $Z$ have the $PEC_{m}[d]$-property, then $X$ has the $PEC_{m}[d]$-property.
We shall prove the theorem via induction on dimension, and henceforth assume that it holds for admissible data $(X^{\prime},Y^{\prime},f^{\prime},U^{\prime},Z^{\prime})$ with $\dim(Y^{\prime})<\dim Y=n$.
6.1. Preliminary reduction
In this subsection, we assume the existence of a non-constant regular function $h:Y\to{\mathbb{A}}^{1}$ such that $Z=g^{-1}(0)$, where $g=h\circ f$. Let $D=f(Z)=h^{-1}(0)$. We prove Theorem 6.1 under these conditions.
@@ 6.1. Preliminary reduction
Let $K\in\operatorname{Perv}(X)$, and let $(K_{U},K_{Z},c,v)$ be the corresponding object in $B_{g}(U,Z)$ (see theorem 2.6). Then
$\chi(K)=\chi(\tilde{j}_{*}K_{U})+\chi(K_{Z})-\chi(\psi_{g}^{u}K_{U})$
Similarly, for $K\in\operatorname{Perv}(Y)$, with the obvious notation we have
$\chi(K)=\chi(j_{*}K_{U})+\chi(K_{D})-\chi(\psi_{h}^{u}K_{U})$
@@
From (2.3.2) and (2.3.1), we obtain
$\begin{split}\chi(K)&=\chi(\ker(v,-\beta))-\chi(\psi_{g}^{u}K_{U})\\ &=\chi(\Xi_{g}K_{U})+\chi(K_{Z})-2\chi(\psi_{g}^{u}K_{U})\\ &=\chi(\tilde{j}_{*}K_{U})+\chi(K_{Z})-\chi(\psi_{g}^{u}K_{U})\end{split}$
The proof of the second part is similar.
@@
    Given an admissible datum $(X,Y,f,U,Z)$ such that $Y$ has the $PEC$-property. If $K_{U}\in\operatorname{Perv}(U)$, then
    $\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})\geq 0.$
@@
    Suppose the $(X,Y,f,U,Z)$ come equipped with compatible ${\mathbb{G}}_{m}$ actions, and that $Y$ satisfies the $PEC_{m}[d]$-property. If $K_{U}\in\operatorname{Perv}_{m}(U)$ then
    $(-1)^{d}(\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U}))\geq 0.$
Proof.
We prove the first part, the proof of the second part is the same. Without loss of generality, we can assume that the ground field $k={\mathbb{C}}$. Given $a\in{\mathbb{N}}$, let $L_{a}$ be pull back to $U$ of the local system on ${\mathbb{C}}^{*}$ with monodromy given by the $a\times a$ Jordan block
We choose (and fix) $a\gg 0$. We may replace $K_{U}$ by $K_{U}\otimes L_{a}$, since the expression $\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})$ just gets multiplied by $\operatorname{rank}L_{a}$. It follows from [24, prop 3.1], that the monodromy of $\psi_{g}^{u}K_{U}$ and $\psi_{h}^{u}K_{U}$ is now trivial. Let $M=j_{!*}K_{U}$. and $\tilde{M}=\tilde{j}_{!*}K_{U}$. By [27, prop 4.7], $\tilde{M}$ corresponds to
$(\tilde{M}_{U}=K_{U},0,0,0)\in B_{g}(U,Z)$
and $M$ corresponds to
$(M_{U}=K_{U},0,0,0)\in B_{h}(U,D$
We have
$j_{*}K_{U}\cong f_{*}\tilde{j}_{*}K_{U}$
by proper base change. Also
$\psi_{h}K_{U}\cong f_{*}\psi_{g}K_{U}$
by [1, exp XIII (2.1.7.1)]. This isomorphism is compatible with monodromy. It follows that this isomorphism restricts to the unipotent parts. Therefore
$\begin{split}\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})&=\chi(j_{*}K_{U})-\chi(\psi_{h}^{u}K_{U})\\ &=\chi(M)\\ &\geq 0\end{split}$
The last inequality follows from the assumptions since $M\in\operatorname{Perv}(Y)$.
@@
Consider an admissible datum $(X,Y,f,U,Z)$ as above. If $Y$ and $Z$ satisfy the $PEC$-property, then so does $X$. In the presence of ${\mathbb{G}}_{m}$-actions, analogous statement holds if $Y$ and $Z$ have the $PEC_{m}[d]$-property.
6.2. Second reduction.
In this subsection, we prove Theorem 6.1 in the setting where we have a function $h:Y\rightarrow{\mathbb{A}}^{1}$ such that $f(Z)\subseteq h^{-1}(0).$ We shall make this assumption for the remainder of this subsection.
@@
Let $Z$ be a Cartier divisor on a variety $X$. Suppose $Z=Z_{1}\cup Z_{2}$ where each $Z_{i}$ is a union of components of $Z$ (with non components in common). If the $PEC$-property holds for each $Z_{i}$, then it also holds for all objects of $Perv(Z)$. The analogous assertion holds in the monodromic setting.
@@
    With notation and hypotheses as in theorem 6.1, suppose furthermore that there exists a function $h:Y\rightarrow{\mathbb{A}}^{1}$ such that $h^{-1}(0)\supseteq f(Z)$. Then the conclusion of the theorem holds.
@@
For (1), let $D_{Y}:=h^{-1}(0)$ and $D_{X}:=g^{-1}(0)$ where $g:=h\circ f:X\rightarrow{\mathbb{A}}^{1}$. Note that $f(D_{X})=f(f^{-1}(D_{Y}))=D_{Y}$. Let $V=Y\setminus D_{Y}\subset U$. It follows that this gives rise to an admissible datum $(X,Y,f,D_{X},V)$. Moreover, we claim that for every perverse sheaf $K\in\operatorname{Perv}(D_{X})$, $\chi(D_{X},K)\geq 0$. In order to see this, first note that the analogous assertion holds for $D_{Y}$ since it holds for $Y$ (and push-forward from a closed subscheme is exact for the perverse t-structure). We may write $D_{X}=D^{\prime}_{X}\cup Z$ as a union of Cartier divisors. Similarly, set $D_{Y}=D^{\prime}_{Y}\cup f(Z)$. Restricting $f$ to $D^{\prime}_{X}$ gives rise to an admissible tuple $(D^{\prime}_{X},D^{\prime}_{Y},f,Z\cap D^{\prime}_{X},U\setminus V)$. We may assume by induction on dimension of $Y$ (or equivalently $X$) that the Euler characteristic of every perverse sheaf on $D^{\prime}_{X}$ is non-negative. Note that this property holds for both $D^{\prime}_{Y}$ and $Z\cap D^{\prime}_{X}$ since it holds for $Y$ and $Z$. Now by Lemma 6.5 above, we may conclude that the Euler characteristic of every perverse sheaf on $D_{X}$ is non-negative, as claimed. Now for the tuple $(X,Y,f,V,D_{X})$, the result follows from Corollary 6.4. Therefore, this proves the desired property for all perverse sheaves on $X$.

### v2

@@ 4. Euler characteristics of perverse sheaves on ${\mathscr{A}}_{g,n}$
5. Key Euler Characteristic Inequality
Given a variety $X$, we say that $X$ has the positive Euler characteristic property, or $PEC$-property, or simply that $X$ is $PEC$, if for every perverse sheaf ${\mathcal{K}}\in\operatorname{Perv}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. If $X$ comes equipped with a ${\mathbb{G}}_{m}$-action, we will use similar terminology in the setting of monodromic perverse sheaves (see 5.2): $X$ has the $PEC_{m}$ property if for all ${\mathcal{K}}\in Perv_{m}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. In the monodromic setting, we shall say that $X$ has the $PEC_{m}[d]$-property if $\chi(X,{\mathcal{K}}[d])\geq 0$ for all ${\mathcal{K}}\in\operatorname{Perv}_{m}(X)$. Clearly $PEC_{m}[d]$ is the same as $PEC_{m}$ if $d$ is even, and $PEC_{m}[-1]$ otherwise.
By an admissible datum, we mean a $5$-tuple $(X,Y,f,U,Z)$, where $f:X\rightarrow Y$ is a (not necessarily proper) birational morphism of irreducible quasiprojective varieties which is an isomorphism over $U\xhookrightarrow{j}Y$, and such that $Z=X\setminus U$ is an (effective) Cartier divisor on X. We say that the datum admits a ${\mathbb{G}}_{m}$-action if ${\mathbb{G}}_{m}$ acts on $X,Y,U,Z$ and $f$ is equivariant with respect to this action. The main result of this section is the following theorem.
Theorem 5.1.
Let $(X,Y,f,U,Z)$ be an admissible datum as above. Suppose that $Y$ and $Z$ have the $PEC$-property. Then $X$ also has the $PEC$-property. The analogous assertion holds in the monodromic setting, i.e. when the datum admits a ${\mathbb{G}}_{m}$-action and $Y$ and $Z$ have the $PEC_{m}[d]$-property, then $X$ has the $PEC_{m}[d]$-property.
Below, we shall prove the theorem via induction on dimension, and henceforth assume that it holds for admissible data $(X^{\prime},Y^{\prime},f^{\prime},U^{\prime},Z^{\prime})$ with $\dim(Y^{\prime})<\dim Y=n$. In the first two subsections, we recall some preliminary material on Belinson’s gluing construction for perverse sheaves, and their monodromic variants. The proof of the main theorem then proceeds via two reductions given in the following two subsections.
5.1. Beilinson’s gluing construction
We recall a basic result of Beilinson [beilinson, morel, reich]. We will follow the notation from [bbd] and write $f_{*}$ etc. for the derived direct image. Also, $\psi_{f},\phi_{f}$ denotes derived nearby/vanishing cycle functors shifted by [1][-1] (so as to preserve perverse sheaves). Let $\psi^{u}_{f}\subset\psi_{f},\phi^{u}_{f}\subset\phi_{f}^{u}$ denote the maximal subfunctors on which the monodromy $T$ acts unipotently. The cancan and varvar maps restrict to these. Let $X$ be a variety with a regular function $h:X\to{\mathbb{A}}^{1}$. Let $Z:=h^{-1}(0)$, U=XZU=X-Z, and let $i:Z\hookrightarrow X,j:U\hookrightarrow X$ denote the canonical inclusions. In this setting, one has the following fundamental result due to Beilinson.
Theorem 5.2 (Beilinson).
$\begin{split}B_{h}(U,Z)&=\{(K_{U},K_{Z},c,v)\mid K_{U}\in\operatorname{Perv}(U),K_{Z}\in\operatorname{Perv}(Z)\\ &c:\psi_{h}^{u}K_{U}\to K_{Z},v:K_{Z}\to\psi_{h}^{u}K_{U},cv=I-T\}\end{split}$
In one direction, $K\in\operatorname{Perv}(X)$ goes to $(j^{*}K,\phi_{h}^{u}K,can,var)$. The inverse is more complicated. Beilinson constructs a “maximal extension” functor $\Xi_{h}:\operatorname{Perv}(U)\to\operatorname{Perv}(X)$ which fits into exact sequences
(5.2.1) $0\to i_{*}\psi^{u}_{h}K_{U}\xrightarrow{\alpha}\Xi_{h}K_{U}\to j_{*}K_{U}\to 0$
$0\to j_{!}K_{U}\to\Xi_{h}K_{U}\xrightarrow{\beta}i_{*}\psi^{u}_{h}K_{U}\to 0$
(5.2.2) $K=H^{0}(i_{*}\psi^{u}_{h}K_{U}\xrightarrow{(\alpha,c)}\Xi_{h}K_{U}\oplus i_{*}K_{Z}\xrightarrow{(\beta,-v)}i_{*}\psi^{u}_{h}K_{U})$
(where this is centered in degree 0).
Remark 5.3.
We observe that the first map in (5.2.2) is injective, and the second is surjective.
Lemma 5.4.
With the above notation, let $K\in\operatorname{Perv}(X)$, and let $(K_{U},K_{Z},c,v)$ be the corresponding object in $B_{h}(U,Z)$. Then
$\chi(K)=\chi(j_{*}K_{U})+\chi(K_{Z})-\chi(\psi_{h}^{u}K_{U})$
Proof.
From (5.2.2), (5.2.1), and the above remark, we obtain
@@ 5.1. Beilinson’s gluing construction
5.2. Monodromic Sheaves
Let $X$ be a variety over an algebraically closed field $k$ such that $char(k)=0$. Suppose that $X$ comes equipped with a ${\mathbb{G}}_{m}$-action denoted by $\theta:{\mathbb{G}}_{m}\times X\rightarrow X.$ For $\lambda\in k^{\times}$, let $\theta_{\lambda}:=\theta(\lambda,-):X\rightarrow X$, and for $x\in X(k)$, let $w_{x}=\theta(-,x):{\mathbb{G}}_{m}\rightarrow X$.
Example 5.5.
A constructible sheaf (respectively complex of sheaves) $\mathcal{K}$} on $X$ is monodromic if $w_{x}^{*}({\mathcal{K}})$ (respectively the cohomology sheaves of $w_{x}^{*}({\mathcal{K}})$) are locally constant for all $x\in X(k)$. We let $Perv_{m}(X)$ denote the full subcategory of monodromic perverse sheaves.
Lemma 5.6.
([Verdier], 3.2) With notation as above, the following are equivalent:
  • $\mathcal{K}$} is monodromic.
Proof.
In the setting of cones and constructible sheaves, this is contained Proposition 3.2 of [Verdier]. The same proof goes through in the current setting. Moreover, the case of complexes can be reduced to sheaves by passing to cohomology sheaves. ∎
Let $X$ be a variety equipped with a ${\mathbb{G}}_{m}$-action, and $h:X\rightarrow{\mathbb{A}}^{1}$ a morphism which is equivariant for the standard ${\mathbb{G}}_{m}$-action on ${\mathbb{A}}^{1}$. Then $Z:=h^{-1}(0)$ and $U=X\setminus Z$ are invariant under ${\mathbb{G}}_{m}$. In this case, we may consider the categories $\operatorname{Perv}_{m}(X),\operatorname{Perv}_{m}(U),$ and $\operatorname{Perv}_{m}(Z)$.
@@ 5.2. Monodromic Sheaves
With notation as above, $\operatorname{Perv}_{m}(X)$ is equivalent to the full subcategory $B_{h,m}(U,Z)\subset B_{h}(U,Z)$ consisting of objects $(K_{U},K_{Z},c,v)$ such that $K_{U}$ and $K_{Z}$ are monodromic.
Proof.
Remark 5.8.
[…]
5.3. Proof of Theorem 5.1: Preliminary reduction
In this subsection, we assume the existence of a non-constant regular function $g:Y\to{\mathbb{A}}^{1}$ such that $Z=h^{-1}(0)$, where $h=g\circ f$. Let $D=f(Z)=g^{-1}(0)$. We prove Theorem 5.1 under these conditions. Below, $\tilde{j}:U\hookrightarrow X$ denotes the natural inclusion.
We begin by recalling a result of Laumon ([laumonEuler]). For a variety, let ${\rm K}_{0}(X)$ denote the Grothendeick group of constructible sheaves. Given a morphism $f:X\rightarrow Y$, we have induced morphisms $Rf_{*},Rf_{!}:{\rm K}_{0}(X)\rightarrow{\rm K}_{0}(Y)$. We also have $f^{-1}:{\rm K}_{0}(Y)\rightarrow{\rm K}_{0}(X)$. For $\mathcal{F}$} in the derived category of constructible sheaves on $X$, $\mathcal{F}$}] denotes its class in ${\rm K}_{0}(X).$
Theorem 5.9.
([laumonEuler]) With notation as above: $Rf_{*}({\mathcal{F}})]=[Rf_{!}({\mathcal{F}})]$ in ${\rm K}_{0}(Y)$ for all ${\mathcal{F}}]\in{\rm K}_{0}(X)$.
We have the following easy consequence which will be used below.
@@
Let $f:X\rightarrow Y$ be a morphism and $D\xhookrightarrow{i}Y$ be a closed subscheme. Let $Z=f^{-1}(D)\xhookrightarrow{\tilde{i}}X$ and $f_{Z}:Z\rightarrow D$ denote the restriction of $f$ to $Z.$ Then for any ${\mathcal{F}}\in{\rm K}_{0}(X)$ we have $i^{-1}Rf_{*}{\mathcal{F}}]=[Rf_{Z,*}(\tilde{i}^{-1}{\mathcal{F}})]\in{\rm K}_{0}(D)$.
Note that if $f$ is proper, then the lemma is an immediate from proper base change.
@@
We have $Rf_{*}({\mathcal{F}})]=[Rf_{!}{\mathcal{F}}]$ in ${\rm K}_{0}(Y)$ by Theorem 5.9. It follows that $i^{-1}Rf_{*}{\mathcal{F}}]=[i^{-1}Rf_{!}{\mathcal{F}}]$ in ${\rm K}_{0}(D)$. By base change for $f_{!}$, we have $i^{-1}Rf_{!}{\mathcal{F}}]=[Rf_{Z,!}(\tilde{i}^{-1}{\mathcal{F}})]$. Another application of Theorem 5.9 gives the desired conclusion. ∎
@@
Let $(X,Y,f,U,Z)$ be an admissible datum such that $Y$ has the $PEC$-property. If $K_{U}\in\operatorname{Perv}(U)$, then
$\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{h}^{u}K_{U})\geq 0.$
Proof.
Without loss of generality, we can assume that the ground field $k={\mathbb{C}}$. Given $a\in{\mathbb{N}}$, let $L_{a}$ be the pull back to $U$ of the local system on $\mathbb{C}$$^{*}$ with monodromy given by the $a\times a$ Jordan block
We choose (and fix) $a\gg 0$. We may replace $K_{U}$ by $K_{U}\otimes L_{a}$, since the expression $\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{h}^{u}K_{U})$ just gets multiplied by $\operatorname{rank}L_{a}$. It follows from [morel, prop 3.1], that the monodromy of $\psi_{g}^{u}K_{U}$ and $\psi_{h}^{u}K_{U}$ is now trivial. Let $M=j_{!*}K_{U}$. and $\tilde{M}=\tilde{j}_{!*}K_{U}$. By [reich, prop 4.7], $\tilde{M}$ corresponds to
$(\tilde{M}_{U}=K_{U},0,0,0)\in B_{h}(U,Z)$
and $M$ corresponds to
$(M_{U}=K_{U},0,0,0)\in B_{g}(U,D).$
Now we note that:
@@
    Since $M\in Perv(Y)$ and $Y$ has PEC, we have $\chi(Y,M)\geq 0$. On the other hand, by Lemma 5.4, this is $\chi(Y,j_{*}K_{U})-\chi(Y,\psi^{u}_{g}(K_{U}))$ since $\phi^{u}_{g}(K_{U})=0$.
@@
    Since $f_{*}(\tilde{j}_{*}K_{U})=j_{*}K_{U}$, $\chi(X,\tilde{j}_{*}K_{U})=\chi(Y,j_{*}K_{U})$.
It remains to show that $\chi(Y,\psi^{u}_{g}(K_{U}))=\chi(X,\psi^{u}_{h}(K_{U}))$. Note that we are in the setting where $\psi^{u}_{g}(K_{U})=i^{-1}j_{*}K_{U}$ and similarly for $\psi_{h}^{u}(K_{U})$ (i.e. vanishing cycles are zero). By Lemma 5.10, we have $Rf_{Z,*}\psi_{h}^{u}(K_{U})]=[Rf_{Z,*}(\tilde{i}^{-1}\tilde{j}_{*}K_{U})]=[i^{-1}Rf_{*}(\tilde{j}_{*}K_{U})]=[i^{-1}j_{*}K_{U}]=[\psi^{u}_{g}(K_{U})]$. ∎
Combining this with Lemma 5.4, we obtain:
@@
Consider an admissible datum $(X,Y,f,U,Z)$ as above and suppose further that there is a function $g:Y\rightarrow{\mathbb{A}}^{1}$ such that $(g\circ f)^{-1}(0)=Z.$ If $Y$ and $Z$ satisfy the $PEC$-property, then so does $X$. In the presence of ${\mathbb{G}}_{m}$-actions, the analogous statement holds if $Y$ and $Z$ have the $PEC_{m}[d]$-property.
5.4. Proof of Theorem 5.1: Second reduction.
In this subsection, we prove Theorem 5.1 in the setting where we have a function $g:Y\rightarrow{\mathbb{A}}^{1}$ such that $f(Z)\subseteq g^{-1}(0).$ We shall make this assumption for the remainder of this subsection.
@@
Let $Z$ be a reduced effective Cartier divisor on a variety $X$. Suppose $Z=Z_{1}\cup Z_{2}$ where each $Z_{i}$ is a union of components of $Z$ (with no components in common). If the $PEC$-property holds for each $Z_{i}$, then it also holds for all objects of $Perv(Z)$. The analogous assertion holds in the monodromic setting.
@@
    With notation and hypotheses as in theorem 5.1, suppose furthermore that there exists a function $g:Y\rightarrow{\mathbb{A}}^{1}$ such that $g^{-1}(0)\supseteq f(Z)$. Then the conclusion of the theorem holds.
@@
For (1), let $D_{Y}:=g^{-1}(0)$ and $D_{X}:=h^{-1}(0)$ where $h:=g\circ f:X\rightarrow{\mathbb{A}}^{1}$. Note that $f(D_{X})=f(f^{-1}(D_{Y}))=D_{Y}$. Let $V=Y\setminus D_{Y}\subset U$. It follows that this gives rise to an admissible datum $(X,Y,f,D_{X},V)$. Moreover, we claim that for every perverse sheaf $K\in\operatorname{Perv}(D_{X})$, $\chi(D_{X},K)\geq 0$. In order to see this, first note that the analogous assertion holds for $D_{Y}$ since it holds for $Y$ (and push-forward from a closed subscheme is exact for the perverse t-structure). We may write $D_{X}=D^{\prime}_{X}\cup Z$ as a union of Cartier divisors. Similarly, set $D_{Y}=D^{\prime}_{Y}\cup f(Z)$. Restricting $f$ to $D^{\prime}_{X}$ gives rise to an admissible tuple $(D^{\prime}_{X},D^{\prime}_{Y},f,Z\cap D^{\prime}_{X},U\setminus V)$. We may assume by induction on dimension of $Y$ (or equivalently $X$) that the Euler characteristic of every perverse sheaf on $D^{\prime}_{X}$ is non-negative. Note that this property holds for both $D^{\prime}_{Y}$ and $Z\cap D^{\prime}_{X}$ since it holds for $Y$ and $Z$. Now by Lemma 5.13 above, we may conclude that the Euler characteristic of every perverse sheaf on $D_{X}$ is non-negative, as claimed. Now for the tuple $(X,Y,f,V,D_{X})$, the result follows from Corollary 5.12. Therefore, this proves the desired property for all perverse sheaves on $X$.